OpenStudy (anonymous):

x=1/25 (22z-9w) please let me know if this correct 1/25(22z-9w)1/25 25x=22z=9w 9w+25x=22z+9w 25x+9w=22z =(25x + 9w)/22 = z

OpenStudy (anonymous):

is this correct!

OpenStudy (anonymous):

u r right :)

OpenStudy (anonymous):

only there r some typing mistakes

OpenStudy (anonymous):

please let me know !

OpenStudy (anonymous):

1/25(22z-9w)=x 25x=22z-9w 9w+25x=22z-9w+9w 25x+9w=22z =(25x + 9w)/22 = z

OpenStudy (anonymous):

thanks!